# Loss Functions and Metrics in Deep Learning. A Review

### v1

@@ 8.1.6 Contrastive Divergence
In many energy-based models, the log-likelihood gradient involves an expectation over all possible data configurations, which is generally intractable to compute exactly. CD approximates this expectation by running a Markov chain for several steps.
@@
where $\theta$ represents the parameters of the model, and $p(x;\theta)$ is the probability of data point $x$, defined as $p(x;\theta)=\frac{e^{-E(x;\theta)}}{Z(\theta)}$, with $E(x;\theta)$ being the energy function and $Z(\theta)$ the partition function.
The gradient of the log-likelihood for the parameters is
$\frac{\partial}{\partial\theta}L(\theta)=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial\theta}\log p(x_{i};\theta)$ (78)
This gradient can be decomposed into two terms: the positive phase, which is easy to compute
$\frac{1}{N}\sum_{i=1}^{N}\frac{\partial E(x_{i};\theta)}{\partial\theta}$ (79)
and the negative phase, which involves an expectation over all model configurations and is generally intractable.
@@
where $\langle\rangle_{p}$ denotes an expectation with respect to the model distribution.
In CD, the negative phase is approximated by running a Markov chain starting from a data point for a few steps and using the resulting sample to estimate the expectation. This leads to the CD-k algorithm, where k is the number of Gibbs sampling steps. The update to the parameters after seeing a data point $x$ is then
$\Delta\theta=\eta(\frac{1}{N}\sum_{i=1}^{N}\frac{\partial E(x_{i};\theta)}{\partial\theta}+\langle\frac{\partial E(x^{\prime};\theta)}{\partial\theta}\rangle_{CD}),$ (81)
where $\eta$ is the learning rate, $\langle\rangle_{CD}$ denotes an expectation to the distribution defined by the Markov chain after k steps, and $x^{\prime}$ is the sample obtained after k steps of Gibbs sampling starting from a data point $x$.
Contrastive Divergence is often more computationally efficient than other methods for training energy-based models, such as Persistent Contrastive Divergence (PCD) [217] or Mean-Field methods [218] because it requires running the Markov chain for a few steps rather than until it reaches equilibrium. However, this introduces a bias in the estimate of the gradient of the log-likelihood. CD is a suitable choice when the bias is acceptable or can be mitigated by running the chain for more steps.
8.2 Image Generation Metrics
@@
8.2.1 Peak Signal-to-Noise Ratio (PSNR)
Peak Signal-to-Noise Ratio (PSNR) is a traditional metric often used to assess the quality of image and video codecs, comparing the original and the reconstructed (compressed and then decompressed) images or videos. It is a measure of the quality of an approximation of an image.
In image generation, PSNR can be used to evaluate the quality of the images generated by models, particularly for tasks like image super-resolution, denoising, inpainting, etc. Here, the PSNR is used to compare the images generated by the model to a ground truth high-quality image.
The PSNR is defined in terms of the mean squared error (MSE), which for two $m\times n$ monochrome images $I$ and $K$ where one of the images is considered a noisy approximation of the other is defined as
@@ 8.2.1 Peak Signal-to-Noise Ratio (PSNR)
where $MAX_{I}$ is the maximum possible pixel value of the image. For instance, for an 8-bit grayscale image, this value would be 255.
PSNR provides an easily interpretable score as expressed in a logarithmic decibel scale. Higher PSNR generally indicates that the reconstruction is of high quality; however, in some cases, perceptually similar images may have low PSNR where simple pixel-wise errors do not well capture structural similarity.
While PSNR could be useful in the context of tasks like super-resolution or denoising where there is a clear ground truth to compare to, it is not particularly effective for tasks like generative image modeling where the goal is to produce new, high-quality images that are not simply reconstructions of existing images [221]. More sophisticated metrics like the Inception Score or Fréchet Inception Distance are typically used for these tasks.
8.2.2 Structural Similarity Index (SSIM)
The Structural Similarity Index (SSIM) [222] can be used in image generation tasks to compare the generated images with the target (real) images. In other words, it measures how similar the synthetic images produced by a generative model are to the actual images. The SSIM index is based on three main factors: luminance, contrast, and structure. The SSIM value ranges from -1 to 1, where 1 indicates perfect similarity, and -1 indicates no similarity.
To calculate the SSIM index, the first step is to compute each image’s mean and standard deviation, then the cross-covariance between the two images, and the product of the standard deviations. The SSIM index is then computed as
@@ 8.2.2 Structural Similarity Index (SSIM)
where $\mu_{x}$ and $\mu_{y}$ are the means of images $x$ and $y$, $\sigma_{x}$ and $\sigma_{y}$ are the standard deviations of images $x$ and $y$, $\sigma_{xy}$ is the cross-covariance between the two images, and $C_{1}$ and $C_{2}$ are constants used to avoid instability.
The SSIM metric is more robust to brightness and contrast changes than other popular image quality metrics, such as the Mean Squared Error (MSE) and the Peak Signal-to-Noise Ratio (PSNR). However, using SSIM alone as an evaluation metric for generative models is insufficient because the ultimate goal of a generative model is not just to reproduce an exact copy of the input image but to understand the underlying data distribution. For this reason, SSIM is often used in conjunction with other metrics, such as Inception Score (IS) and Fréchet Inception Distance (FID), to evaluate the performance of generative models more comprehensively.
8.2.3 Inception Score (IS)
The Inception Score (IS) quantifies the quality and diversity of generated images [219]. It is computed by combining two scores: the marginal likelihood of the generated images and their quality or diversity. The marginal likelihood is calculated by feeding the generated images into an Inception-v3 classifier trained on the ImageNet dataset and taking the average of the softmax probabilities. Their overall scores determine the quality of the images, while the entropy of these scores reflects the diversity.
The formula for the inception score can be expressed as
$IS=e^{(\mathbb{E}_{x\sim p_{g}(x)}[KL(p(y|x)||p(y))])},$ (85)
where $p_{g}(x)$ is the distribution of generated images, $p(y|x)$ is the class conditional probability for each image and class, and $p(y)$ is the marginal likelihood of each class.
The Inception Score provides a single number that reflects the quality of generated images and the diversity of these objects and it is easy to calculate using pre-trained Inception models that don’t require the true data distribution as input.
However, the Inception Score has some limitations that make it insufficient to comprehensively understand the model performance:
@@ 8.2.3 Inception Score (IS)
    Inception Model Dependence: Relies heavily on the Inception model, which is pre-trained on the ImageNet dataset. This may limit its usefulness in domains that are very different from ImageNet, as the features learned by the Inception model might not be relevant for those domains.
@@
    Mismatch with Human Perception: It does not always align with human perception of image quality. Images can achieve high IS by simply being diverse and recognizable, even if they don’t look real.
@@
    Lack of Discrimination Between Modes: It can be high for models that cover many modes of the true data distribution but fail to accurately capture the frequency of those modes.
@@
    Doesn’t Measure Mode Collapse: As mentioned before, a problem in training GANs is mode collapse, where the model generates a limited range of images. The Inception Score can’t effectively detect mode collapse as long as the few modes that the GAN does generate are diverse enough.
@@
    Unreliable for Low Number of Samples: When computed over a small number of samples, the Inception Score can become unreliable due to high variance.
8.2.4 Fréchet Inception Distance (FID)
Unlike the Inception Score, the Fréchet Inception Distance (FID) [220] considers the statistics of the images generated by the model and compares them to the statistics of real images. It measures the distance between these two distributions in a feature space provided by a specific layer of the Inception network.
Mathematically, the FID is defined as follows. Let us denote the generated images as $X$ and the real images as $Y$. Each dataset is passed through an Inception network, and the activations at a specific layer are extracted. The resulting activations are denoted as $X^{\prime}$ and $Y^{\prime}$ respectively. Assuming that $X^{\prime}$ and $Y^{\prime}$ are multivariate Gaussian distributions with means $\mu_{x}$ and $\mu_{y}$, and covariance matrices $\Sigma_{x}$ and $\Sigma_{y}$, respectively, the Fréchet distance between these two multivariate Gaussian distributions is defined as
@@ 8.2.4 Fréchet Inception Distance (FID)
where $Tr$ is the trace of a matrix, which is the sum of its diagonal elements, $(\Sigma_{x}\Sigma_{y})^{1/2}$ is the matrix square root of the product of the covariance matrices, which is well-defined since the covariance matrices are positive semi-definite.
FID has the following desirable properties:
@@
    Lower value is better: The lower the FID score, the closer the generated image distribution is to the real image distribution.
@@
    Considers both mean and covariance: FID considers both the means and the covariance of the feature representations, giving it an advantage over metrics considering only one of these aspects.

### v3

@@ 4.5.1 Image Generation Loss functions
Contrastive Divergence (CD) is a method used in training Energy-Based Models (EBMs) to approximate the gradient of the log-likelihood, which involves an expectation over all possible data configurations. Calculating this expectation exactly is typically intractable. CD approximates it by running a Markov chain for a limited number of steps.
@@
where $\theta$ represents the parameters of the model, and $p(x;\theta)$ is the probability of a data point $x$, defined as $p(x;\theta)=\frac{e^{-E(x;\theta)}}{Z(\theta)}$, with $E(x;\theta)$ being the energy function and $Z(\theta)$ the partition function.
The gradient of the log-likelihood with respect to the parameters is:
$\frac{\partial}{\partial\theta}L(\theta)=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial\theta}\log p(x_{i};\theta).$ (102)
This gradient can be decomposed into two terms:
1. Positive Phase: This term is computed from the data distribution and is straightforward to compute:
$\frac{1}{N}\sum_{i=1}^{N}\frac{\partial E(x_{i};\theta)}{\partial\theta}.$ (103)
2. Negative Phase: This term involves an expectation over the model’s distribution and is generally intractable:
@@ 1. Positive Phase: This term is computed from the data distribution and is straightforward to compute:
where $\langle\cdot\rangle_{p}$ denotes an expectation with respect to the model distribution.
In CD, the negative phase is approximated by running a Markov chain starting from a data point for a few steps, obtaining a sample to estimate the expectation. This leads to the CD-k algorithm, where $k$ represents the number of Gibbs sampling steps. The update to the parameters after observing a data point $x$ is then:
$\Delta\theta=\eta\left(\frac{1}{N}\sum_{i=1}^{N}\frac{\partial E(x_{i};\theta)}{\partial\theta}-\langle\frac{\partial E(x^{\prime};\theta)}{\partial\theta}\rangle_{CD}\right),$ (105)
where $\eta$ is the learning rate, $\langle\cdot\rangle_{CD}$ denotes the expectation with respect to the distribution defined by the Markov chain after $k$ steps, and $x^{\prime}$ is the sample obtained after $k$ steps of Gibbs sampling starting from the data point $x$.
Contrastive Divergence is often more computationally efficient than other methods, such as Persistent Contrastive Divergence (PCD) [245] or Mean-Field methods [246], because it requires running the Markov chain for only a few steps instead of until reaching equilibrium. However, this introduces a bias in the gradient estimate of the log-likelihood. CD is suitable when the bias is acceptable or can be mitigated by increasing the number of steps in the Markov chain.
4.5.2 Image Generation Metrics
@@
Peak Signal-to-Noise Ratio (PSNR)
Peak Signal-to-Noise Ratio (PSNR) is a traditional metric used to quantify the quality of image and video codecs by comparing the original and reconstructed (compressed and then decompressed) images or videos. It measures the fidelity of the reconstruction, indicating how closely the reconstructed image resembles the original.
In image generation tasks, PSNR is often used to evaluate the quality of the generated images, particularly in applications such as image super-resolution, denoising, and inpainting. PSNR compares the generated image to a ground truth high-quality image, providing an objective assessment of reconstruction quality.
The PSNR is defined in terms of the mean squared error (MSE), which measures the average squared differences between the original and the reconstructed images. For two $m\times n$ monochrome images $I$ and $K$, where one image is a noisy approximation of the other, MSE is calculated as:
@@
where $MAX_{I}$ is the maximum possible pixel value of the image. For example, in an 8-bit grayscale image, this value is 255.
PSNR is expressed in decibels (dB) and provides an easily interpretable score: higher PSNR values generally indicate better quality reconstructions. However, it is important to note that PSNR may not always correlate with perceived visual quality. For example, two images with similar perceptual quality can have significantly different PSNR values if they differ in pixel-wise detail.
While PSNR is useful for evaluating tasks with a clear ground truth, such as super-resolution or denoising, it is less effective for generative tasks where the goal is to produce new, high-quality images rather than exact reconstructions. In such cases, metrics like the Inception Score (IS) or Fréchet Inception Distance (FID) are often preferred, as they better capture perceptual quality and diversity [249].
Structural Similarity Index (SSIM)
The Structural Similarity Index (SSIM) [250] is a metric used to assess the similarity between two images. It is particularly useful in image generation tasks for comparing generated images with target (real) images, providing a quantitative measure of how closely synthetic images produced by a generative model resemble actual images.
SSIM considers three main components: luminance, contrast, and structure. These components reflect perceptual characteristics that are important in human vision, making SSIM a more perceptually relevant measure compared to simpler metrics like Mean Squared Error (MSE) or Peak Signal-to-Noise Ratio (PSNR). The SSIM value ranges from -1 to 1, where 1 indicates perfect similarity, 0 indicates no correlation, and -1 indicates complete dissimilarity.
The calculation of the SSIM index involves the following steps:
  1. 1.
    Luminance Comparison: Measures the similarity in brightness between two images.
  2. 2.
    Contrast Comparison: Evaluates the similarity in contrast or intensity variation.
  3. 3.
    Structural Comparison: Assesses the similarity in structural information, focusing on the pattern of pixel intensities.
The formula for SSIM between two images $x$ and $y$ is given by:
@@
where:
  • $\mu_{x}$ and $\mu_{y}$ are the means of the pixel intensities in images $x$ and $y$, representing luminance.
  • $\sigma_{x}$ and $\sigma_{y}$ are the standard deviations of the pixel intensities in $x$ and $y$, representing contrast.
  • $\sigma_{xy}$ is the covariance of the pixel intensities between $x$ and $y$, representing structural similarity.
  • $C_{1}$ and $C_{2}$ are small constants added to avoid instability when the denominator is close to zero, typically defined as $C_{1}=(K_{1}L)^{2}$ and $C_{2}=(K_{2}L)^{2}$, where $L$ is the dynamic range of the pixel values (e.g., 255 for 8-bit grayscale images), and $K_{1}$ and $K_{2}$ are small constants.
SSIM is more robust to variations in brightness and contrast compared to MSE and PSNR because it considers local patterns in pixel intensities that are consistent with human visual perception. However, while SSIM is valuable for evaluating the perceptual quality of images, it does not fully capture the diversity or the distributional properties of the generated data.
For a comprehensive evaluation of generative models, SSIM is often used alongside other metrics such as the Inception Score (IS) and Fréchet Inception Distance (FID). These additional metrics help assess the model’s ability to capture the underlying data distribution and generate diverse and high-quality images.
Inception Score (IS)
The Inception Score (IS) is a metric used to evaluate the quality and diversity of images generated by generative models, such as GANs. The score was introduced by Salimans et al. [247] and is computed using an Inception-v3 classifier pre-trained on the ImageNet dataset.
The IS combines two key aspects: the quality of the generated images and the diversity of the objects within them. The score is based on the idea that a good generative model should produce images that are both recognizable (high quality) and varied across different classes (diversity).
To calculate the Inception Score, the generated images are passed through the Inception-v3 network, and the resulting softmax probabilities are used. The score is calculated using the following formula:
$IS=\exp(\mathbb{E}_{x\sim p_{g}(x)}[KL(p(y|x)||p(y))]),$ (109)
where:
  • $p_{g}(x)$ represents the distribution of the generated images.
  • $p(y|x)$ is the conditional probability of class $y$ given image $x$, as predicted by the Inception model.
  • $p(y)=\mathbb{E}_{x\sim p_{g}(x)}[p(y|x)]$ is the marginal class distribution over all generated images.
The Kullback-Leibler (KL) divergence term measures how different the conditional distribution $p(y|x)$ is from the marginal distribution $p(y)$. A high KL divergence indicates that the model generates images with distinct class labels (diversity), while a low entropy in $p(y|x)$ suggests high confidence in the predictions (quality).
The Inception Score provides a single scalar value that reflects both the quality and diversity of the generated images. It is easy to calculate and does not require access to the true data distribution, making it a convenient evaluation metric.
The limitations of the Inception Score include the following:
@@
    Dependence on the Inception Model: IS relies heavily on the Inception-v3 model, which is trained on ImageNet. This dependence may limit the score’s applicability in domains significantly different from ImageNet, as the model’s learned features might not generalize well.
@@
    Mismatch with Human Perception: The IS does not always correlate with human judgments of image quality. It can give high scores to images that are diverse and recognizable, even if they lack photorealism.
@@
    Inability to Detect Mode Collapse: The score may not effectively detect mode collapse (where the model generates a limited variety of images) if the few modes present are sufficiently diverse.
@@
    Insensitivity to Class Balance: The IS does not account for the frequency with which different classes are generated, potentially overlooking scenarios where the model produces certain classes disproportionately.
@@
    Unreliable with Few Samples: When calculated with a small number of samples, the IS can be unreliable due to high variance, making it less robust in such cases.
Fréchet Inception Distance (FID)
The Fréchet Inception Distance (FID) [248] is a metric used to evaluate the quality of images generated by a model by comparing their statistical properties to those of real images. Unlike the Inception Score, which primarily focuses on the class distributions of generated images, FID measures the distance between the distributions of real and generated images in a feature space, typically using the activations of a specific layer in an Inception network.
To compute FID, generated images (denoted as $X$) and real images (denoted as $Y$) are passed through an Inception network, and activations from a chosen layer are extracted. These activations are represented as $X^{\prime}$ and $Y^{\prime}$, respectively. The FID assumes that these activations follow a multivariate Gaussian distribution, characterized by the mean and covariance matrix. Let $\mu_{x}$ and $\mu_{y}$ denote the means of the activations for $X^{\prime}$ and $Y^{\prime}$, and $\Sigma_{x}$ and $\Sigma_{y}$ their respective covariance matrices. The Fréchet distance between these two Gaussian distributions is defined as:
@@
where:
  • $||\mu_{x}-\mu_{y}||^{2}$ is the squared Euclidean distance between the means of the two distributions, measuring differences in their central tendency.
  • $Tr$ denotes the trace of a matrix, which is the sum of its diagonal elements.
  • $(\Sigma_{x}\Sigma_{y})^{1/2}$ is the matrix square root of the product of the covariance matrices, representing the similarity in the spread and orientation of the distributions. This term is well-defined since covariance matrices are positive semi-definite.
The FID has two key advantages. It considers both the average features and the diversity of the generated images by taking into account the mean and covariance of the feature representations. Additionally, FID is less affected by noise compared to the Inception Score, making it a more dependable measure of generative model quality.
However, the FID has limitations. Similar to the Inception Score, FID relies on the Inception network, which is pre-trained on the ImageNet dataset. This reliance can limit its effectiveness when evaluating images from domains significantly different from ImageNet, as the extracted features may not be as relevant. Additionally, FID assumes that the feature distributions are Gaussian, which may not always hold true, potentially affecting the accuracy of the metric.
FID has become a standard metric for evaluating generative models because it effectively captures differences in both quality and diversity between the real and generated images.
5 Natural Language Processing
Natural Language Processing (NLP) is a subfield of artificial intelligence that focuses on the interaction between computers and human languages. It involves the development of algorithms and models that allow computers to understand, interpret, and generate human language in a way that is both meaningful and useful. NLP encompasses a wide range of tasks, from basic text processing and understanding to complex language generation and translation.
The main goal of NLP is to bridge the gap between human communication and machine understanding, allowing more natural and efficient interactions with technology. This includes enabling machines to comprehend the subtleties of human language, such as syntax, semantics, context, and sentiment.
Tha main tasks in NLP are the following:
@@ 5 Natural Language Processing
    Text Classification: Involves categorizing text into predefined classes or categories. Examples include sentiment analysis, spam detection, and topic classification.
@@
    Language Modeling: Predicts the probability of a sequence of words. It is a foundational task in NLP, crucial for applications like text generation and autocomplete.